# CangLing-KnowFlow: A Unified Knowledge-and-Flow-fused Agent for Comprehensive Remote Sensing Applications

### v1

@@ 5 School of Information and Communication Technology, Griffith University, Nathan, QLD 4111, Australia
The automated and intelligent processing of massive remote sensing (RS) datasets is critical in Earth observation (EO). Existing automated systems are normally task-specific, lacking a unified framework to manage diverse, end-to-end workflows–from data preprocessing to advanced interpretation–across diverse RS applications. To address this gap, this paper introduces CangLing-KnowFlow, a unified intelligent agent framework that integrates a Procedural Knowledge Base (PKB), Dynamic Workflow Adjustment, and an Evolutionary Memory Module. The PKB, comprising 1,008 expert-validated workflow cases across 162 practical RS tasks, guides planning and substantially reduces hallucinations common in general-purpose agents. During runtime failures, the Dynamic Workflow Adjustment autonomously diagnoses and replans recovery strategies, while the Evolutionary Memory Module continuously learns from these events, iteratively enhancing the agent’s knowledge and performance. This synergy enables CangLing-KnowFlow to adapt, learn, and operate reliably across diverse, complex tasks. We evaluated CangLing-KnowFlow on the KnowFlow-Bench, a novel benchmark of 324 workflows inspired by real-world applications, testing its performance across 13 top Large Language Model (LLM) backbones, from open-source to commercial. Across all complex tasks, CangLing-KnowFlow surpassed the Reflexion baseline by at least 4% in Task Success Rate. As the first most comprehensive validation along this emerging field, this research demonstrates the great potential of CangLing-KnowFlow as a robust, efficient, and scalable automated solution for complex EO challenges by leveraging expert knowledge (Knowledge) into adaptive and verifiable procedures (Flow).
@@ 1 Introduction
Recent advances in Earth observation technologies have led to a rapid growth in the availability of remote sensing (RS) data, marked by increasing volume, diversity, and temporal resolution [vivone2024deep, zhang2022progress]. While this data surge offers transformative opportunities for applications such as environmental monitoring and disaster response, it also highlights the limitations of traditional RS analysis pipelines. These conventional methods–largely dependent on manually crafted workflows designed by domain experts–struggle with inefficiency, high labor demands, and limited scalability, creating a critical bottleneck to timely and automated extraction of actionable intelligence from large and complex datasets [zhang2019remotely, zhang2025core].
In parallel, breakthroughs in Artificial Intelligence (AI)–driven by the rise of LLMs–have introduced the AI Agent as a transformative operational paradigm. AI agents possess the ability to interpret high-level tasks, plan and execute multi-step workflows, dynamically invoke specialized tools, and perform self-correction. These capabilities enable the seamless integration of previously disjointed, human-dependent processes into coherent, automated pipelines. This paradigm shift from human-driven methodologies to machine-autonomous operations opens a new pathway toward realizing the next generation of autonomous Earth observation systems [li2025segearth].
Building a robust and reliable RS agent is challenging due to the inherent limitations of general-purpose agent architectures. First, monolithic single-agent systems often produce erroneous or scientifically unsound plans–a phenomenon known as planning hallucination [zhu2025knowagent]–especially when dealing with long-horizon tasks requiring deep domain expertise. Second, while traditional Multi-Agent Systems promote collaboration, their decentralized and open-ended communication protocols can introduce excessive overhead and goal misalignment, undermining the procedural determinism and result reproducibility crucial for scientific analysis. Finally, fully fixed workflows, although stable and compliant, lack the flexibility to accommodate the variability inherent in RS data (e.g., sensor types, weather conditions, temporal dynamics) and to recover from unexpected runtime failures, yielding a persistent trade-off between reliability and adaptability [shi2025flowagent].
@@
Despite their considerable potential, the application of agent-based systems in RS remains in an exploratory phase, characterized by two primary technical trajectories. The first encompasses tool-augmented agents, including single-agent systems for tool invocation [xu2024rs, wei2025geotool, chen2024llm, ning2024llm] and multi-agent frameworks tailored for specific data types [lin2025shapefilegpt, lee2025multi, yu2024mineagent]. While these efforts validate the feasibility of agents interfacing with specialized tools, their operational scope is often confined to single-turn interactions or predefined toolsets, exhibiting limited capacity for long-term, autonomous planning and adaptation. The second trajectory is defined by unified Vision Language Models (VLMs) [hu2025rsgpt, zhan2025skyeyegpt, soni2025earthdial], which attempt to handle various RS tasks end-to-end within a single massive model. Although these models demonstrate notable multitasking capabilities, their monolithic nature precludes the precise orchestration of complex external toolchains, and their lack of workflow-level robustness renders them difficult to diagnose and correct upon failure. The collective architectural deficiencies of these approaches define a significant research gap. Figure 1 provides a conceptual illustration of this gap by contrasting the limitations of existing paradigms (Figure 1a) with our proposed CangLing-KnowFlow framework, which features an integrated design by leveraging expert knowledge into adaptive working flows (Figure 1b).
To bridge the divergent technical paths in Figure 1 by balancing the flexibility of tool use with the rigor of scientific processes, the Agentic Workflow has emerged as a more advanced organizational paradigm[li2025designing, feng2025earth]. This paradigm leverages sophisticated orchestration to guide agents through a structured, collaborative process, thereby ensuring analytical transparency while retaining the ability to dynamically compose toolchains. Initial explorations in the geospatial domain have begun to adopt this approach [bhattaram2025geoflow]. However, the preliminary implementations suffer from notable limitations. First, they rely excessively on the generalized capabilities of LLMs for planning. Although LLMs possess extensive general knowledge, their outputs, optimized for universality, often lack the requisite specificity and depth for specialized scientific domains. An operational agent must be infused with the tacit knowledge, operational heuristics, and industry standards of domain experts, a capability underdeveloped in existing workflows. Second, they predominantly follow a static plan-once, execute-sequentially model, rendering them brittle in the face of runtime failures caused by data or environmental factors. Third, current workflow agents typically operate without a persistent memory mechanism and, therefore, cannot learn from past successes or failures as some generalized agents [shinn2023reflexion], limiting their potential for continuous intelligence growth.
@@
    Autonomous real-time adaptation ensures robust performance in dynamic conditions. Our framework empowers the agent with real-time monitoring and dynamic adjustment capabilities. When a procedural step fails, the agent can autonomously diagnose the issue and replan the subsequent path, demonstrating strong robustness in unpredictable scenarios.
@@
    A persistent memory mechanism drives continuous, experience-driven learning. We introduce a long-term memory system that enables the agent to learn from operational experience. This system solidifies successful adjustments into new workflows and attributes failures to heuristic knowledge, addressing the critical learning gap in current agentic workflows.
@@
The remainder of this paper details the architectural design of our framework and the implementation of the above three core mechanisms, followed by experimental setup, results analysis, and a discussion of the potential to advance automation and intelligence in RS science.
@@ 2.1 General-Purpose LLM Agents
The conceptual foundation of our research lies in the rapid evolution of general-purpose agents powered by LLMs. Early advances demonstrated that LLMs could perform rudimentary planning via techniques such as Chain-of-Thought prompting. A pivotal development was the introduction of frameworks like ReAct [yao2023react], which synergized reasoning and acting by interleaving thought processes with tool-based actions, enabling agents to interact with external environments to solve tasks. This led to a wave of research into enhancing agent capabilities, particularly in generalized tool use, supported by the creation of comprehensive benchmarks and datasets [li2023api, tang2023toolalpaca]. Building upon this, subsequent research has focused on enhancing agent autonomy and intelligence. For example, some approaches have incorporated self-reflection and memory mechanisms to allow agents to learn from past failures [shinn2023reflexion], while others have integrated tree-search algorithms to improve planning and decision-making capabilities [zhou2024language]. Concurrently, multi-agent conversational frameworks have been proposed to tackle complex problems through collaborative task decomposition [wu2024autogen]. Despite these advances, a common limitation of general-purpose agents is their reliance on the intrinsic, generalized knowledge of the underlying LLM. When applied to specialized scientific domains, this often leads to planning hallucination, where the generated plans lack scientific rigor and procedural validity [zhu2025knowagent], a challenge highlighted extensively by existing evaluation benchmarks [liu2023agentbench].
@@ 2.2 Agentic Workflows
To address the limitations of unstructured, general-purpose agents, the paradigm of the Agentic Workflow has emerged. This approach imposes a structured, often procedural, constraint on agent behavior to increase reliability and ensure compliance with domain-specific constraints. A key focus within this paradigm is the infusion of external knowledge to guide the planning process. Some studies have proposed augmenting agents with explicit knowledge bases to constrain the generation of action plans, thereby improving their factuality and relevance [zhu2025knowagent, ou2025automind]. Others have explored the inherent tension between procedural compliance and operational flexibility, developing frameworks that allow agents to adhere to a predefined workflow while also managing unexpected, out-of-scope situations [shi2025flowagent]. This relates directly to the need for dynamic adjustment in real-world applications. More recently, the concept of memory has been integrated into workflows, enabling agents to learn and reuse successful procedural patterns from past experiences [wang2024agent]. Some advanced systems even attempt to automate the generation of the workflow itself [zhang2024aflow]. However, the systematic definition, orchestration, dynamic adaptation, and experiential optimization of such workflows for highly specialized domains like RS remain significant open research questions.
@@ 2.3 AI Agents in Remote Sensing and Geoscience
The application of AI agents to RS and geoscience is a nascent, but rapidly developing field. Early efforts have focused predominantly on tool augmentation, demonstrating that agents can interpret user commands to invoke specific geospatial analysis tools or models. These include single-agent systems for tool execution [xu2024rs, wei2025geotool, zhang2024geogpt, akinboyewa2025gis, chen2024llm, ning2024llm] and multi-agent frameworks for specialized data processing or collaborative analysis [lin2025shapefilegpt, lee2025multi, tang2025geosr, pantiukhin2025accelerating, xie2025rag, yu2024mineagent]. While foundational, these systems are often limited to simple, single-turn tool calls and lack the capability to orchestrate complex, multi-step analytical workflows.
More recently, GeoFlow [bhattaram2025geoflow] introduced an agentic workflow paradigm to the geospatial domain, representing a significant step forward. However, this and similar initial implementations exhibit three critical deficiencies that our research aims to address. First, they rely on the LLM’s general knowledge for planning and lack a mechanism for injecting codified, expert-validated procedural knowledge, making them prone to domain-specific errors. Second, their workflows are largely static, planned once at the beginning of a task, and lack the robustness to dynamically adjust to runtime failures. Third, they are memoryless, precluding the ability to learn from or evolve through accumulated task logs and retrieved historical cases. Furthermore, while benchmarks for evaluating tool-use or specific spatial cognition abilities are emerging [shabbir2025thinkgeo], a comprehensive benchmark designed to rigorously assess an agent’s advanced capabilities in complex workflow orchestration, dynamic adaptation, and experiential learning is still absent. This paper introduces the CangLing-KnowFlow framework and corresponding benchmark to fill these critical gaps.
@@ 3.1 Framework Overview
The CangLing-KnowFlow framework is engineered to address the core challenge of fusing deep domain knowledge with a flexible, adaptive procedural flow for automating complex RS tasks. The central design principle is to create an agent that not only executes predefined plans, but which also adapts dynamically to unforeseen circumstances and learns systematically from its operational experience. As shown in Figure 2, the architecture of our proposed CangLing-KnowFlow is composed of three primary, synergistic components, integrated under a central Orchestrator Agent. This agent serves as the cognitive core, responsible for interpreting user requests, planning the execution, and initiating interventions. It draws plans from an expert-informed Procedural Knowledge Base (PKB), a repository of codified, expert-validated RS workflows. The concrete execution is handled by a Dynamic Execution Engine, which interfaces with the toolset and provides real-time feedback on the outcome of each action. Finally, an Evolutionary Memory Module records all execution traces, enabling the agent to learn from experience by analyzing successes and failures to refine its knowledge over time.
@@
The decision-making process of the Orchestrator Agent can be formally modeled as a Markov Decision Process (MDP) [shi2025flowagent]. At each time step $t$, the agent selects an action $a_{t}$ based on the history of the current task and guidance from its knowledge sources. We extend the traditional workflow agent formulation to incorporate explicitly the influence of long-term, cross-task experience. The action $a_{t}$ is determined by the agent’s policy $\mathcal{A}$ as follows:
@@
The practical implementation of the above formal model is realized through a structured operational cycle: (1) Task Decomposition and Workflow Retrieval, where the Orchestrator parses the user’s goal and retrieves the most relevant workflow template from the PKB (see Section 3.2). (2) Situated Planning, where the Orchestrator adapts the retrieved template to the specific data and contextual parameters of the current task. (3) Monitored Execution, where the Dynamic Execution Engine executes the planned steps sequentially, with the Orchestrator supervising the process. (4) Dynamic Adjustment, which is triggered upon execution failure, prompting the Orchestrator to diagnose the issue and re-plan the workflow. (5) Experiential Learning, which occurs post-task, where the Evolutionary Memory Module analyzes the complete execution trace to update the system’s knowledge base. This operational cycle is formally detailed in Algorithm 1. The subsequent sections provide a detailed exposition of the PKB, the dynamic adjustment mechanism, and the evolutionary memory module.

### v2

@@ 5 School of Information and Communication Technology, Griffith University, Nathan, QLD 4111, Australia
Automated and intelligent processing of massive remote sensing datasets is critical in Earth observation. Existing automated systems are normally task-specific, lacking a unified framework to manage diverse, end-to-end workflows from data preprocessing to advanced interpretation across diverse RS applications. To fill this gap, this paper introduces CangLing-KnowFlow, a unified intelligent agent framework that integrates a procedural knowledge base, a dynamic workflow adjustment mechanism, and an evolutionary memory module. The procedural knowledge base, comprising 1,008 expert-validated workflow cases across 162 practical remote sensing tasks, guides planning and substantially reduces hallucinations common in general-purpose agents. During runtime failures, the dynamic workflow adjustment module autonomously diagnoses and replans recovery strategies, while the evolutionary memory module continuously learns from these events, iteratively enhancing the agent’s knowledge and performance. This synergy enables CangLing-KnowFlow to adapt, learn, and operate reliably across diverse and complex tasks. We evaluated CangLing-KnowFlow on the KnowFlow-Bench, a novel benchmark of 324 workflows inspired by real-world applications, testing its performance across 13 top large language model backbones, from open-source to commercial. Across all complex tasks, CangLing-KnowFlow surpassed the Reflexion baseline by at least 4% in task success rate. As the first most comprehensive validation along this emerging field, this research demonstrates the great potential of CangLing-KnowFlow as a robust, efficient, and scalable automated solution for complex Earth observation challenges by leveraging expert knowledge (Know) into adaptive and verifiable procedures (Flow).
@@ 1 Introduction
Recent advances in Earth observation technologies have led to a rapid growth in the availability of remote sensing (RS) data, marked by increasing volume, diversity, and temporal resolution [vivone2024deep, zhang2022progress]. While this data surge offers transformative opportunities for applications such as environmental monitoring and disaster response, it also highlights the limitations of traditional RS analysis pipelines. These conventional methods largely dependent on manually crafted workflows designed by domain experts, therefore struggling with inefficiency, high labor demands, and limited scalability, which creates a critical bottleneck to timely and automated extraction of actionable intelligence from large and complex datasets [zhang2019remotely, zhang2025core].
In parallel, breakthroughs in Artificial Intelligence (AI), driven by the rise of large language models (LLMs), have introduced the AI Agent as a transformative operational paradigm. AI agents possess the ability to interpret high-level tasks, plan and execute multi-step workflows, dynamically invoke specialized tools, and perform self-correction. These capabilities enable the seamless integration of previously disjointed, human-dependent processes into coherent and automated pipelines. This paradigm shift from human-driven methodologies to machine-autonomous operations opens a new pathway toward realizing the next generation of autonomous Earth observation systems [li2025segearth].
Building a robust and reliable RS agent is challenging due to the inherent limitations of general-purpose agent architectures. First, monolithic single-agent systems often produce erroneous or scientifically unsound plans, which is a phenomenon known as planning hallucination [zhu2025knowagent], especially when dealing with long-horizon tasks requiring deep domain expertise. Second, although traditional multi-agent systems promote collaboration, their decentralized communication protocols can introduce excessive overhead and misaligned goals. This could create uncertainty in the process and compromise the reproducibility required for scientific analysis. Finally, fully fixed workflows, although stable and compliant, lack the flexibility to accommodate the variability inherent in RS data (e.g., sensor types, weather conditions, temporal dynamics) and recover from unexpected runtime failures, yielding a persistent trade-off between reliability and adaptability [shi2025flowagent].
@@
Despite their considerable potential, the application of agent-based systems in RS remains in an exploratory phase, currently following two main directions. The first encompasses tool-augmented agents, including single-agent systems for tool invocation [xu2024rs, wei2025geotool, chen2024llm, ning2024llm] and multi-agent frameworks tailored for specific data types [lin2025shapefilegpt, lee2025multi, yu2024mineagent]. While these efforts validate the feasibility of agents interfacing with specialized tools, their operational scope is often confined to single-turn interactions or predefined toolsets, exhibiting limited capacity for long-term, autonomous planning and adaptation. The second direction is defined by unified Vision Language Models (VLMs) [hu2025rsgpt, zhan2025skyeyegpt, soni2025earthdial], attempting to handle various RS tasks end-to-end within a single massive model. Although these models demonstrate notable multitasking capabilities, on one hand, their monolithic nature precludes the precise orchestration of complex external toolchains, and on the other hand, their lack of workflow-level robustness renders them difficult to diagnose and correct upon failure. The architectural deficiencies of the existing paradigms mentioned above constitute a significant research gap, which is conceptually illustrated in Figure 1 (a).
More recently, agentic workflows have emerged as a solution to bridge these divergent technical paths, blancing tool flexibility with the strict requirements of scientific workflows [li2025designing, feng2025earth]. By orchestrating agents through structured processes, this paradigm ensures analytical transparency while retaining the ability to dynamically compose toolchains. Although early geospatial works have explored this direction [bhattaram2025geoflow], it still suffers from notable limitations. First, they rely excessively on the generalizaion capabilities of LLMs for planning. Although LLMs possess extensive general knowledge, their outputs, optimized for universality, often lack the requisite specificity and depth for specialized scientific domains. An operational agent must be infused with the tacit knowledge, operational heuristics, and industry standards of domain experts, a capability underdeveloped in existing workflows. Second, they predominantly follow a static plan-once, execute-sequentially model, making them fragile when faced with runtime failures caused by data or environmental factors. Third, current workflow agents typically operate without a persistent memory mechanism and, therefore, cannot learn from past successes or failures as some generalized agents [shinn2023reflexion], limiting their potential for continuous intelligence growth.
@@
    Autonomous real-time adaptation ensures robust performance in dynamic conditions. Our framework empowers the agent with real-time monitoring and dynamic adjustment capabilities. When a procedural step fails, the agent can autonomously diagnose the issue and replan the subsequent path, demonstrating strong robustness in unseen scenarios.
@@
    A persistent memory mechanism drives continuous, experience-driven learning. We introduce a long-term memory system that enables the agent to learn from operational experience. This solidifies successful adjustments into new workflows and attributes failures to heuristic knowledge, addressing the critical learning gap in current agentic workflows.
@@
The remainder of this paper is structured as follows. Section 2 provides a comprehensive review of the literature, covering general LLM-based agents, agentic workflows, and AI agents for Earth observation tasks. Section 3 introduces the proposed Cangling-KnowFlow framework and elaborates on its three key components. Section 4 outlines the constructed toolset database, the curated KnowFlow-Bench dataset, and the evaluation protocols. Section 5 reports the experimental settings and results, followed by an in-depth analysis. Finally, Section 6 concludes this paper and discusses future research directions.
@@ 2.1 General-Purpose LLM Agents
The conceptual foundation of our research lies in the rapid evolution of general-purpose agents powered by LLMs. Early efforts demonstrated that LLMs could perform preliminary planning via techniques such as Chain-of-Thought prompting. A pivotal development was the introduction of frameworks like ReAct [yao2023react], which synergized reasoning and acting by interleaving thought processes with tool-based actions, enabling agents to interact with external environments to solve tasks. It further led to a wave of research into enhancing agent capabilities, particularly in generalized tool use, supported by the creation of comprehensive benchmarks and datasets [li2023api, tang2023toolalpaca]. Building upon this, subsequent research has focused on enhancing agent autonomy and intelligence. For example, some approaches have incorporated self-reflection and memory mechanisms to allow agents to learn from past failures [shinn2023reflexion], while others have integrated tree-search algorithms to improve planning and decision-making capabilities [zhou2024language]. Concurrently, multi-agent conversational frameworks have been proposed to tackle complex problems through collaborative task decomposition [wu2024autogen]. Despite these advances, a common limitation of general-purpose agents is their reliance on the intrinsic, generalized knowledge of the underlying LLM. In specialized domains, this often leads to ’planning hallucinations’. As highlighted by recent benchmarks [liu2023agentbench], these agents tend to generate plans that appear plausible but are scientifically incorrect.
@@ 2.2 Agentic Workflows
To address the limitations of unstructured, general-purpose agents, the paradigm of the Agentic Workflow has emerged. This approach imposes a structured, often procedural, constraint on agent behavior to increase reliability and ensure compliance with domain-specific constraints. A key focus within this paradigm is the infusion of external knowledge to guide the planning process. Some studies have proposed augmenting agents with explicit knowledge bases to constrain the generation of action plans, thereby improving their factuality and relevance [zhu2025knowagent, ou2025automind]. Others have explored the inherent tension between procedural compliance and operational flexibility, developing frameworks that allow agents to adhere to a predefined workflow while also managing unexpected, out-of-scope situations [shi2025flowagent]. This relates directly to the need for dynamic adjustment in real-world applications. More recently, the concept of memory has been integrated into workflows, enabling agents to learn and reuse successful procedural patterns from past experiences [wang2024agent]. Some advanced systems even attempt to automate the generation of the workflow itself [zhang2024aflow]. However, the systematic definition, orchestration, dynamic adaptation, and experiential optimization of such workflows for highly specialized domains like RS remain open questions.
@@ 2.3 AI Agents in Remote Sensing and Geoscience
The application of AI agents to RS and Geoscience is a nascent but rapidly developing field. Early efforts have focused predominantly on tool augmentation, demonstrating that agents can interpret user commands to invoke specific geospatial analysis tools or models. These include single-agent systems for tool execution [xu2024rs, wei2025geotool, zhang2024geogpt, akinboyewa2025gis, chen2024llm, ning2024llm] and multi-agent frameworks for specialized data processing or collaborative analysis [lin2025shapefilegpt, lee2025multi, tang2025geosr, pantiukhin2025accelerating, xie2025rag, yu2024mineagent]. While foundational, these systems are often limited to simple, single-turn tool calls and lack the capability to orchestrate complex, multi-step analytical workflows.
More recently, GeoFlow [bhattaram2025geoflow] introduced the Agentic Workflow paradigm to the geospatial domain, representing a significant step forward. However, this and similar early attempts exhibit three critical deficiencies that are worth noting. First, they rely on the LLM’s general knowledge for planning and lack a mechanism for injecting codified, expert-validated procedural knowledge, making them prone to domain-specific errors. Second, their workflows are largely static, planned once at the beginning of a task, and lack the robustness to dynamically adjust to runtime failures. Third, they are memoryless, precluding the ability to learn from or evolve through accumulated task logs and retrieved historical cases. Furthermore, while benchmarks for evaluating tool-use or specific spatial cognition abilities are emerging [shabbir2025thinkgeo], a comprehensive benchmark designed to rigorously assess an agent’s advanced capabilities in complex workflow orchestration, dynamic adaptation, and empirical learning is still absent, which constitutes the motivation of this paper.
@@ 3.1 Framework Overview
The CangLing-KnowFlow framework is developed to fill the above-mentioned critical gaps by addressing the core challenge of fusing deep domain knowledge with a flexible, adaptive procedural flow for automating complex RS tasks. The key design principle is to create an agent that not only executes predefined plans, but which also adapts dynamically to unforeseen circumstances and learns systematically from its operational experience. As shown in Figure 2, the architecture of our proposed CangLing-KnowFlow is composed of three primary, synergistic components, integrated under a central Orchestrator Agent. This agent serves as the cognitive core, responsible for interpreting user requests, planning the execution, and initiating interventions. It draws plans from an expert-informed Procedural Knowledge Base (PKB), a repository of codified, expert-validated RS workflows. The concrete execution is handled by a Dynamic Execution Engine, which interfaces with the toolset and provides real-time feedback on the outcome of each action. Finally, an Evolutionary Memory Module records all execution traces, enabling the agent to learn from experience by analyzing successes and failures to refine its knowledge over time.
@@
The decision-making process of the Orchestrator Agent can be formally modeled as a Markov Decision Process (MDP) [shi2025flowagent]. At each time step $t$, the agent selects an action $a_{t}$ based on the history of the current task and guidance from its knowledge sources. We extend the traditional workflow agent formulation to incorporate the influence of long-term, cross-task experience explicitly. The action $a_{t}$ is determined by the agent’s policy $\mathcal{A}$ as follows:
@@
The practical implementation of the above formal model is realized through a structured operational cycle: (1) Task Decomposition and Workflow Retrieval, where the Orchestrator parses the goal of users and retrieves the most relevant workflow template from the PKB (see Section 3.2). (2) Situated Planning, where the Orchestrator adapts the retrieved template to the specific data and contextual parameters of the current task. (3) Monitored Execution, where the Dynamic Execution Engine executes the planned steps sequentially, with the Orchestrator supervising the process. (4) Dynamic Adjustment, which is triggered upon execution failure, prompting the Orchestrator to diagnose the issue and re-plan the workflow. (5) Empirical Learning, which occurs post-task, where the Evolutionary Memory Module analyzes the complete execution trace to update the system’s knowledge base. This operational cycle is formally detailed in Algorithm 1. The subsequent sections will provide a detailed discussion of the three components, i.e., the PKB, the dynamic adjustment mechanism, and the evolutionary memory module.